OpenStudy (anonymous):

Identify the period for the trigonometric function: f(t)= 3cot(pi*t) A.pi/3 B.3 C.1 D. pi

OpenStudy (anonymous):

2pi/pi

OpenStudy (jhannybean):

Take for example: \(g(x)) = \sin\left(\pi x+\dfrac{\pi}{2}\right)+3\) It follows the same order, except the format now would be \(y=A\sin(Bx)\) therefpre the period is: \(\dfrac{2\pi}{\pi} = 2\)

OpenStudy (anonymous):

doesn't it just say identify so would it be just D. pi?

OpenStudy (jhannybean):

Yeah, period of a regular cotangent function is pi.

OpenStudy (anonymous):

Can I ask another question? its about the same thing...

OpenStudy (jhannybean):

Wait a minute...

OpenStudy (jhannybean):

I found my mistake

OpenStudy (jhannybean):

Period for tangent/cotangent: \(P = \dfrac{\pi}{|B|}\) Sine and cosine functions have a period of \(2\pi\) whereas cotangent, and tangent functions have a period of \(\pi\)

OpenStudy (anonymous):

so its c?

OpenStudy (jhannybean):

Therefore, \(P = \dfrac{\pi}{\pi} = 1~\checkmark\)

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

No problem, relearned something again today, haha.